Solve the system of equations.
x + 3y + 5z = -4
3y + 3z = -6
z = 2

Answer by mananth(16949)

x + 3y + 5z = -4
3y + 3z = -6
z = 2
3y + 3z = -6
z = 2
3y+3*2=-6
3y+6=-6
3y=-12
/3
y=-4
x+3y+5z=-4
x+3(-4)+5(2)=-4
x-12+10=-4
x-2=-4
x=-2